The work done when two mole of an ideal gas is compressed from a volume of 5 m3 to 1 dm3 at 300 K, under a pressure of 100 kPa is
1
Expert's answer
2017-05-09T10:26:09-0400
The work done on an ideal gas is
W=p(V_1-V_2 )=100000(5-0.001)=499900 J=499.9 kJ.